[figure: 9.7 cm]

TRANSPORT PROPERTIES OF THE U_{1−x}Th_{x}Pt SYSTEM

V.H. Tran and R. Troć W. Trzebiatowski Institute of Low Temperature and Structure Research, Polish Academy of Sciences, P.O.Box 1410, 50-950 Wrocław, Poland

UPt was reported to crystallize in two distinct crystal structures with the orthorhombic CrB-[1] or the monoclinic PdBi-type [2]. Two magnetic transitions into a ferromagnetic state at 27 K [3] and 19 K [4] have also merged. We present the measurements of electrical resistivity (ρ(T)) and magnetoresistance (MR) on the solid solutions U_{1−x}Th_{x}Pt, in which the parent compound UPt orders ferromagnetically below T_{C} = 21 K. The following points are noteworthy about our experimental results: (i) ρ(T) of all the samples show typical spin-fluctuation behaviour with the AT^{2}-dependence in the ordered state. (ii) In agreement with magnetization and specific heat measurements [5], we observe a continuous suppression of the ferromagnetic ordering upon the Th-substitution.(iii) The MR of UPt is negative at low temperatures, reaching a high value of -22 % at T_{C} and B = 8 T. (iv) The value of MR increases with x and becomes positive for Th-rich compositions. We discuss these results in the framework of existing models.

[1] A.E. Dwight, in: Developments in the Structural Chemistry of Alloy Phases, ed. B.C. Giessen (Plenum, New York 1969), p. 181.

[2] A. Dommann and F. Hulliger, Solid State Comun. 65 (1988) 1093.

[3] J.G. Huber, M.B. Maple and D. Wohlleben, J. Magn. Magn. Mater. 1 (1975) 58.

[4] J.J.M. Franse et al. Physica B 123 (1983) 11.

[5] R. Troć at al. (to be published).

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : V.H. Tran

Address for correspondence : W.Trzebiatowski Institute of Low Temperature and Structure Research, Polish Academy of Sciences, P.O.Box 1410, 50-950 Wrocław, Poland

Email address : vhtran@int.pan.wroc.pl